# Harmonic functions on Tutte’s embeddings and linearized Monge–Ampère equation

### v1

@@ 4.1. Oscillations estimate via the Dirichlet energy
In this section we recall a well-known equicontinuity estimate for discrete harmonic functions via their Dirichlet energies; see Lemma 4.3 and Remark 4.2 below. In its simplest form, this estimate dates back to Lusternik [Lusternik1926]; see also [skopenkov-Dirichlet, gurel-gurevich-jerison-nachmias] and references therein.
@@
Given two subsets $A,B\subset U_{\delta}$, we denote by $\mathrm{EL}_{U}^{\delta}(A\leftrightarrow B)$ the (edge) extremal length of paths connecting $A$ and $B$ inside $U_{\delta}$; this quantity is also known as the effective resistance between $A$ and $B$ in the electrical network $U_{\delta}$. Similarly, given $A^{\ast},B^{\ast}\subset U_{\delta}^{\ast}$, we denote by $\mathrm{EL}_{U}^{\delta,\ast}(A^{\ast}\leftrightarrow B^{\ast})$ the extremal length of paths connecting $A^{\ast}$ and $B^{\ast}$ inside $U_{\delta}^{\ast}$. We address the reader to [lyons-peres-book, Chapter 2] or to [chelkak-robust, Section 6][binder2024orthodiagonal, Section 2.2] and references therein for the definition and basic properties of the discrete extremal length.
@@
where we also used the fact that the Dirichlet energies of the harmonic function $H$ and of its harmonic conjugate $H^{\ast}$ are equal. We can now bound $\mathrm{EL}^{\delta,\ast}_{A(p,r,2r)}(\gamma_{M}^{\ast}\leftrightarrow\gamma_{m}^{\ast})$ from above by the extremal length of circuits separating two boundary components of the annulus. Classically (see [ford1956maximal, Theorem 1]), the latter equals the inverse of the extremal length $\mathrm{EL}^{\delta}_{A(p,r,2r)}(\partial B(p,r)\leftrightarrow\partial B(p,2r))$, where $\partial B(p,r)$ denotes the set of vertices in $A(p,r,2r)_{\delta}$ incident to vertices from $\overline{B}(p,r)_{\delta}$, and the same for $\partial B(p,2r)$. Therefore,
@@ 4.3. Proofs of Theorem 2 and Theorem 3 on the assumption $\varphi\in{\mathrm{C}}^{3}$
The main convergence results of our paper, Theorem 2 and Theorem 3 claim the convergence of discrete harmonic functions $H_{\delta}$ to solutions of the equation ${\mathcal{L}}_{\varphi}h=0$ under no local ‘non-degeneracy’ assumptions on the embeddings $\Gamma_{\delta}$. At the same time, in applications of such results to the dimer model one often needs to prove convergence of the gradients of $H_{\delta}$. (For instance, see [berestycki-lis-qian-free-dimers, Section 1.7] for a short discussion of the link between dimer model observables in this setup and discrete Green’s functions.) In this section we strengthen Theorem 2 and Theorem 3 to the convergence of gradients of $H_{\delta}$ under a mild additional regularity assumption (EXP-FAT) on $\Gamma_{\delta}$ that originated in [CLR1]. We say that a face of $\Gamma_{\delta}$ is $\rho$-fat if it can be triangulated so that each triangle contains a disc of radius $\rho$.
Assumption (EXP-FAT) (cf. [CLR1, Assumption 5.9]).
@@
The framework developed in [CLR1] provides a ‘discrete version’ of the computations made above when a smooth potential $\varphi$ is replaced with a piecewise linear potential $\Phi_{\delta}$ and $h$ with a discrete harmonic function $H_{\delta}$ on $\Gamma_{\delta}$. Splitting each face of $\Gamma_{\delta}$ into triangles and extending $H_{\delta}$ inside these triangles linearly we can view $H_{\delta}$ as a piecewise linear function and its gradient $F_{\delta}:=\partial_{w}H_{\delta}$ as a piecewise constant function. (Note the advantage of the discrete setup: $F_{\delta}$ is well-defined pointwise for each discrete potential $\Phi_{\delta}$.) The discrete analog of the 1-form (4.6) is then the 1-form $F_{\delta}d{\mathcal{T}}_{\delta}-\overline{F}_{\delta}d{\mathcal{O}}_{\delta}$ on the t-surface $\Theta_{\delta}$ associated with $\Gamma_{\delta}$. (See [CLR1, Section 4.2 and Section 5] for an extension of this 1-form to the whole t-surface from black faces.) As above, this gives a certain discrete harmonicity property for each of the functions $\operatorname{\mathrm{Re}}[\alpha F_{\delta}]$, $\alpha\in{\mathbb{T}}$, in the coordinate ${\mathcal{T}}_{\delta}+\alpha^{2}{\mathcal{O}}_{\delta}$; see [CLR1, Section 4.3] for details. Importantly, if $\Gamma_{\delta}$ has property (CONV)/(LIP), with no additional assumptions on $\Gamma_{\delta}$, then the corresponding random walks on T-graphs ${\mathcal{T}}_{\delta}+\alpha^{2}{\mathcal{O}}_{\delta}$ are uniformly elliptic starting from the scale $\delta$. This leads to the following:
@@
Note that the gradient of $H_{\delta}$ can be seen as a t-black-holomorphic function on the t-embedding ${\mathcal{T}}_{\delta}=\frac{1}{2}({\mathcal{T}}_{\delta}+{\mathcal{O}}_{\delta})$; see the discussion at the end of Section 2.4 and [CLR1, Proposition 8.2(ii)]. Recall also that the mapping from $\Gamma_{\delta}$ to ${\mathcal{T}}_{\delta}$ is bi-Lipschitz on convex sets starting from the scale $\delta$. Item (i) follows from [CLR1, Proposition 4.17 and Proposition 5.7]. Items (ii) and (iii) follow from [CLR1, Proposition 6.13 and Theorem 6.17]. ∎
@@
Let us emphasize that in Theorem 4.8 we do not impose any additional regularity assumption on the potential $\varphi\in{\mathrm{C}}^{1,1}$ besides the uniform convexity (1.4). In particular, one can use Theorem 4.8 to prove that $h\in{\mathrm{C}}^{1,\beta}_{\mathrm{loc}}$ for each (weak) solution $h\in W^{1,2}_{\mathrm{loc}}$ of the equation ${\mathcal{L}}_{\varphi}h=0$ with uniformly convex potential $\varphi$ by discretizing the operator ${\mathcal{L}}_{\varphi}$; see [Gutierrez-Truyen-gradient2011] for related results.
@@
and denote by $\widetilde{P}_{\delta}$ the upper convex hull of the set $\{(v_{\delta};\Phi_{\delta}(v_{\delta})),v_{\delta}\in V_{\delta}\}\subset{\mathbb{C}}\times{\mathbb{R}}$. Since $\Phi_{\delta}$ is convex, $\widetilde{P}_{\delta}$ is the supergraph of a convex function $\widetilde{\Phi}_{\delta}$ that is defined on the convex hull of $V_{\delta}$ and is linear on faces of a certain graph whose vertices belong to $V_{\delta}$. We denote this graph by $\widetilde{\Gamma}_{\delta}$ and define the weight of an edge of $\widetilde{\Gamma}_{\delta}$ as the jump of the gradient of $\widetilde{\Phi}_{\delta}$ between two adjacent faces divided by the length of this edge. (In other words, the weighted graph $\widetilde{\Gamma}_{\delta}$ and the modified potential $\widetilde{\Phi}_{\delta}$ are related to each other in the usual way; e.g., see (1.1).) It follows from (1.4) that for each small enough $\varepsilon>0$ there exists $\delta_{0}(\varepsilon)>0$ such that for all $\delta\leq\delta_{0}(\varepsilon)$
@@
The approximation trick used above does not allow us to conclude that the gradients $\nabla H_{\delta}$ are always uniformly bounded. However, it allows one to prove that $H_{\delta}$ are ‘differentiable at a mesoscopic scale’; the result recently obtained in [pechersky-rate2025] for harmonic functions on orthodiagonal tilings. To this end, note that in the arguments given above one can set $\widetilde{r}=R-\delta^{\alpha}$, $\alpha<1$, and $\varepsilon=C_{0}\delta$ with a big enough constant $C_{0}$ depending on constants in property (CONV) only. If one replaces $\sqrt{\varepsilon}$ in the proof of (4.8) with $\delta^{\frac{1}{2}(1+\alpha)}$, the right-hand side of this estimate reads as $O(\delta^{\frac{1}{2}\beta(1-\alpha)})$. Thus,
@@ 5. Equivalence of properties (CONV)/(LIP) and (RW)
Denote $Z=\sum_{v\in B(w,r)}\mu_{\delta}(v)$ and $\nu=Z^{-1}\mu_{\delta}1\!\!1_{B(w,r)}$. Recall that $Z\geq cr^{2}$ as we assume property (RW). Let $X_{t}^{\nu}$ be the random walk on $\Gamma_{\delta}$ stopped upon leaving $B(w,r)$ with $X_{0}$ distributed according to $\nu$. Provided that $C$ is large enough, Proposition 2.7 and a standard large deviation estimate (e.g., see [CLR1, Proposition 6.1]) imply that for each $\alpha\in{\mathbb{T}}$ one has
@@
As $z\mapsto{\mathcal{O}}_{\delta}(z)$ is a $1$-Lipschitz function, the mapping $D\ni z\mapsto z-\alpha^{2}{\mathcal{O}}_{\delta}(z)$ has no overlaps; see also [CLR1, proof of Proposition 4.3]. Let $Q$ be the $2C\delta$-neighborhood of the segment $z_{0};z_{1}]$. It follows from the last estimate that

### v2

@@ 4.1. Oscillations estimate via the Dirichlet energy
In this section we recall a well-known equicontinuity estimate for discrete harmonic functions via their Dirichlet energies; see Lemma 4.3 and Remark 4.2 below. In its simplest form, this estimate dates back to Lusternik [33]; see also [42, 23] and references therein.
@@
Given two subsets $A,B\subset U_{\delta}$, we denote by $\mathrm{EL}_{U}^{\delta}(A\leftrightarrow B)$ the (edge) extremal length of paths connecting $A$ and $B$ inside $U_{\delta}$; this quantity is also known as the effective resistance between $A$ and $B$ in the electrical network $U_{\delta}$. Similarly, given $A^{\ast},B^{\ast}\subset U_{\delta}^{\ast}$, we denote by $\mathrm{EL}_{U}^{\delta,\ast}(A^{\ast}\leftrightarrow B^{\ast})$ the extremal length of paths connecting $A^{\ast}$ and $B^{\ast}$ inside $U_{\delta}^{\ast}$. We address the reader to [32, Chapter 2] or to [15, Section 6][5, Section 2.2] and references therein for the definition and basic properties of the discrete extremal length.
@@
where we also used the fact that the Dirichlet energies of the harmonic function $H$ and of its harmonic conjugate $H^{\ast}$ are equal. We can now bound $\mathrm{EL}^{\delta,\ast}_{A(p,r,2r)}(\gamma_{M}^{\ast}\leftrightarrow\gamma_{m}^{\ast})$ from above by the extremal length of circuits separating two boundary components of the annulus. Classically (see [21, Theorem 1]), the latter equals the inverse of the extremal length $\mathrm{EL}^{\delta}_{A(p,r,2r)}(\partial B(p,r)\leftrightarrow\partial B(p,2r))$, where $\partial B(p,r)$ denotes the set of vertices in $A(p,r,2r)_{\delta}$ incident to vertices from $\overline{B}(p,r)_{\delta}$, and the same for $\partial B(p,2r)$. Therefore,
@@ 4.3. Proofs of Theorem 2 and Theorem 3 on the assumption $\varphi\in{\mathrm{C}}^{3}$
The main convergence results of our paper, Theorem 2 and Theorem 3 claim the convergence of discrete harmonic functions $H_{\delta}$ to solutions of the equation ${\mathcal{L}}_{\varphi}h=0$ under no local ‘non-degeneracy’ assumptions on the embeddings $\Gamma_{\delta}$. At the same time, in applications of such results to the dimer model one often needs to prove convergence of the gradients of $H_{\delta}$. (For instance, see [2, Section 1.7] for a short discussion of the link between dimer model observables in this setup and discrete Green’s functions.) In this section we strengthen Theorem 2 and Theorem 3 to the convergence of gradients of $H_{\delta}$ under a mild additional regularity assumption (EXP-FAT) on $\Gamma_{\delta}$ that originated in [11]. We say that a face of $\Gamma_{\delta}$ is $\rho$-fat if it can be triangulated so that each triangle contains a disc of radius $\rho$.
Assumption (EXP-FAT) (cf. [11, Assumption 5.9]).
@@
The framework developed in [11] provides a ‘discrete version’ of the computations made above when a smooth potential $\varphi$ is replaced with the Maxwell–Cremona potential $\Phi_{\delta}$ and $h$ with a discrete harmonic function $H_{\delta}$ on $\Gamma_{\delta}$. Splitting each face of $\Gamma_{\delta}$ into triangles and extending $H_{\delta}$ inside these triangles linearly we can view $H_{\delta}$ as a piecewise linear function and its gradient $F_{\delta}:=\partial_{w}H_{\delta}$ as a piecewise constant function. (Note the advantage of the discrete setup: $F_{\delta}$ is well-defined pointwise for each discrete potential $\Phi_{\delta}$.) The discrete analog of the 1-form (4.6) is then the 1-form $F_{\delta}d{\mathcal{T}}_{\delta}-\overline{F}_{\delta}d{\mathcal{O}}_{\delta}$ on the t-surface $\Theta_{\delta}$ associated with $\Gamma_{\delta}$. (See [11, Section 4.2 and Section 5] for an extension of this 1-form to the whole t-surface from black faces.) As above, this gives a certain discrete harmonicity property for each of the functions $\operatorname{\mathrm{Re}}[\alpha F_{\delta}]$, $\alpha\in{\mathbb{T}}$, in the coordinate ${\mathcal{T}}_{\delta}+\alpha^{2}{\mathcal{O}}_{\delta}$; see [11, Section 4.3] for details. Importantly, if $\Gamma_{\delta}$ has property (CONV)/(LIP), with no additional assumptions on $\Gamma_{\delta}$, then the corresponding random walks on T-graphs ${\mathcal{T}}_{\delta}+\alpha^{2}{\mathcal{O}}_{\delta}$ are uniformly elliptic starting from the scale $\delta$. This leads to the following:
@@
Note that the gradient of $H_{\delta}$ can be seen as a t-black-holomorphic function on the t-embedding ${\mathcal{T}}_{\delta}=\frac{1}{2}({\mathcal{T}}_{\delta}+{\mathcal{O}}_{\delta})$; see the discussion at the end of Section 2.4 and [11, Proposition 8.2(ii)]. Recall also that the mapping from $\Gamma_{\delta}$ to ${\mathcal{T}}_{\delta}$ is bi-Lipschitz on convex sets starting from the scale $\delta$. Item (i) follows from [11, Proposition 4.17 and Proposition 5.7]. Items (ii) and (iii) follow from [11, Proposition 6.13 and Theorem 6.17]. ∎
@@
Let us emphasize that in Theorem 4.8 we do not impose any additional regularity assumption on the potential $\varphi\in{\mathrm{C}}^{1,1}$ besides the uniform convexity (1.4). In particular, one can use Theorem 4.8 to prove that $h\in{\mathrm{C}}^{1,\beta}_{\mathrm{loc}}$ for each (weak) solution $h\in W^{1,2}_{\mathrm{loc}}$ of the equation ${\mathcal{L}}_{\varphi}h=0$ with uniformly convex potential $\varphi$ by discretizing the operator ${\mathcal{L}}_{\varphi}$; see [24] for related results.
@@
and denote by $\widetilde{P}_{\delta}$ the upper convex hull of the set $\{(v_{\delta};\Phi_{\delta}(v_{\delta})),v_{\delta}\in\widetilde{V}_{\delta}\}\subset{\mathbb{C}}\times{\mathbb{R}}$. Since $\Phi_{\delta}$ is convex, $\widetilde{P}_{\delta}$ is the supergraph of a convex function $\widetilde{\Phi}_{\delta}$ that is defined on the convex hull of $\widetilde{V}_{\delta}$ and is linear on faces of a certain graph whose vertices belong to $\widetilde{V}_{\delta}$. We denote this graph by $\widetilde{\Gamma}_{\delta}$ and define the weights on edges of $\widetilde{\Gamma}_{\delta}$ using Maxwell–Cremona correspondence (In other words, the weighted graph $\widetilde{\Gamma}_{\delta}$ and the modified potential $\widetilde{\Phi}_{\delta}$ are related to each other in the usual way; e.g., see (1.1).) Note that the weights are positive since $\widetilde{\Phi}_{\delta}$ is convex. It follows from (1.4) that for each small enough $\varepsilon>0$ there exists $\delta_{0}(\varepsilon)>0$ such that for all $\delta\leq\delta_{0}(\varepsilon)$
@@
The approximation trick used above does not allow us to conclude that the gradients $\nabla H_{\delta}$ are always uniformly bounded. However, it allows one to prove that $H_{\delta}$ are ‘differentiable at a mesoscopic scale’; the result recently obtained in [38] for harmonic functions on orthodiagonal tilings. To this end, note that in the arguments given above one can set $\widetilde{r}=R-\delta^{\alpha}$, $\alpha<1$, and $\varepsilon=C_{0}\delta$ with a big enough constant $C_{0}$ depending on constants in property (CONV) only. If one replaces $\sqrt{\varepsilon}$ in the proof of (4.8) with $\delta^{\frac{1}{2}(1+\alpha)}$, the right-hand side of this estimate reads as $O(\delta^{\frac{1}{2}\beta(1-\alpha)})$. Thus,
@@ 5. Equivalence of properties (CONV)/(LIP) and (RW)
Denote $Z=\sum_{v\in B(w,r)}\mu_{\delta}(v)$ and $\nu=Z^{-1}\mu_{\delta}1\!\!1_{B(w,r)}$. Recall that $Z\geq cr^{2}$ as we assume property (RW). Let $X_{t}^{\nu}$ be the random walk on $\Gamma_{\delta}$ stopped upon leaving $B(w,r)$ with $X_{0}$ distributed according to $\nu$. Provided that $C$ is large enough, Proposition 2.7 and a standard large deviation estimate (e.g., see [11, Proposition 6.1]) imply that for each $\alpha\in{\mathbb{T}}$ one has
@@
As $z\mapsto{\mathcal{O}}_{\delta}(z)$ is a $1$-Lipschitz function, the mapping $D\ni z\mapsto z-\alpha^{2}{\mathcal{O}}_{\delta}(z)$ has no overlaps; see also [11, proof of Proposition 4.3]. Let $Q$ be the $2C\delta$-neighborhood of the segment $z_{0};z_{1}]$. It follows from the last estimate that